# From planar to annular to toroidal bracket polynomials for pseudo knots and links

### v1

@@ 3.1. Preliminaries
Figure 16. The lift of a toroidal pseudo link to a pseudo link in $T^{2}\times I$ (cf. [5]).
@@
Theorem 3.2.
[5] Two oriented ${\rm H}$-mixed pseudo links in $S^{3}$ are isotopic if and only if any two oriented ${\rm H}$-mixed pseudo link diagrams of theirs differ by planar isotopies, a finite sequence of the classical and the pseudo Reidemeister moves for the moving parts of the mixed pseudo links (recall Figure 2), and moves that involve the fixed and the moving parts, called mixed Reidemeister moves, comprising the moves MR2, MR3, and MPR3, as exemplified in Figure 9 and the mixed R3 moves, as exemplified in Figure 18.
@@ 3.2. The toroidal pseudo bracket polynomial
A $(p,q)$-torus knot, where $p$ and $q$ are relatively prime integers, is an essential simple closed curve in the torus winding $p$ times around the meridian and $q$ times along the longitude of the torus, with slope $p/q$. Similarly, a $(p^{\prime},q^{\prime})$-torus link or a $(kp,kq)$-torus link on $k$ components is a set of $k$ parallel copies of a $(p,q)$-torus knot, where $p^{\prime}=k\cdot p$ and $q^{\prime}=k\cdot q$. Torus knots and links are characterized by their slopes, so a $(p,q)$-torus knot is the same curve as a $(-p,-q)$-torus knot while a $(p,-q)$-torus knot is the same curve as a $(-p,q)$-torus knot. Otherwise, any two torus knots or links with different slopes are non-isotopic.
@@
A $(p,q)$-torus knot, considered as a classical knot, is the same as a $(q,p)$-torus knot, since the two solid tori making up the 3-sphere, and bearing the knot on their common boundary, can interchange their roles. However, in the knot theory of the thickened torus, these two knots are distinct since the host manifold is fixed.
All torus knots and links with distinct slopes along with the standard null-homotopic unknot constitute initial conditions for the pseudo bracket polynomial of toroidal pseudo link diagrams. So, we assign a specific variable $s_{p,q}$ to each $(p,q)$-torus knot and we define the following:
@@
Let $K$ be an oriented toroidal pseudo link diagram. The toroidal pseudo bracket polynomial of $K$, denoted $\langle K\rangle_{T^{2}}$ or simply $\langle K\rangle$, is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,s_{p,q}\right]$, for $p$ and $q$ pairs of coprime integers with $q$ non negative, defined inductively by means of the following rules, where a punctured plane in the 5th rule represents the torus:
@@
Let $K$ be an oriented toroidal pseudo link. The reduced toroidal pseudo bracket polynomial of $K$, denoted $\langle K\rangle^{r}_{T^{2}}$ or simply $\langle K\rangle^{r}$, is a 5-variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,x^{\pm 1},y\right]$, defined inductively by means of the same rules as for the polynomial $\langle K\rangle_{T^{2}}(A,V,H,s_{p,q})$, except for the fifth rule, in which $s_{p,q}$ is replaced by $x^{p}y^{q}$, for $p$ and $q$ pairs of coprime integers with $q$ non negative.
Note that the last rule can only be applied to a toroidal state diagram that contains only a standard null-homotopic unknot. We also point out that two torus knot components of a toroidal pseudo link diagram, which have different slopes, will necessarily form crossings, which will be smoothed by the skein relations of Definition 3.4 (and Definition 3.5 for the reduced case). So, a state diagram of the toroidal pseudo bracket polynomial can only contain a single torus knot or link. Therefore, together with the comment above Definition 3.4 we have that the toroidal pseudo bracket polynomial and the reduced one are well defined.
@@
The toroidal pseudo bracket polynomial and the reduced toroidal pseudo bracket polynomial are invariant under regular isotopy of oriented toroidal pseudo links.
@@
Figure 19. Inclusion relations (cf. [5]): a disc in the torus and an annulus in the torus.
@@
Figure 20. Inclusion relation of the thickened torus in a 3-ball (cf. [5]).
@@
Figure 21. Inclusion relation of the thickened torus in the solid torus (cf. [5]).
@@
(e) Consider a torus link consisting of two different torus knot components, say a $(1,0)$-torus knot and a $(0,1)$-torus knot which intersect in a precrossing. Assigning orientations to the two components we observe that reversing the orientation of one component will result in different bracket polynomials, due to the fact that the roles of $V$ and $H$ in the third rule of the toroidal pseudo bracket are not interchangeable. The situation is illustrated in Figure 23. The left-hand side toroidal pseudo link, $L_{1}$, is assigned the polynomial:
@@
The pseudo trefoil on the top right side of the figure, denoted $K_{r}$, is not of annular type, so when projecting on the annulus, a virtual crossing will appear, as illustrated in Figure 24 bottom right, which is neither a real crossing nor a precrossing (cf. [13]). The use of virtual crossings in this context comes in handy (only) for illustration purposes, as discussed in [5], in order to represent toroidal diagrams in the annulus or in the once punctured disk. We stress that the isotopy moves of virtual knot theory do not apply here. The skein tree analysis of $K_{r}$ is depicted in Figure 25.
@@ 3.3. The ${\rm H}$-mixed pseudo bracket polynomial
We now consider the $(kp,kq)$-torus links, which form states for the toroidal pseudo bracket polynomial of Definition 3.4, which here correspond to ${\rm H}$-mixed links. Figure 26 illustrates a $(3,2)$-torus knot represented in two ways as an ${\rm H}$-mixed link. The left-hand illustration of the figure uses the common representation of ${\rm H}$, while in the right-hand illustration ${\rm H}$ is represented in closed braid form. As we observe, in the left-hand illustration appear three moving crossings, which should be smoothened according to the bracket rules, a situation we can avoid using, for example, the representation of the $(3,2)$-torus knot in the right-hand illustration.
@@
We are now ready to define:
@@
Let ${\rm H}\cup K$ be an oriented ${\rm H}$-mixed pseudo link diagram, with ${\rm H}$ being represented in closed braid form. The ${\rm H}$-mixed pseudo bracket polynomial of ${\rm H}\cup K$ is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,s_{p,q}\right]$, defined by means of the same inductive rules as the ones for the toroidal mixed pseudo bracket polynomial, except for the $(kp,kq)$-torus link diagram in the fifth rule of Definition 3.4, which is substituted by the corresponding ${\rm H}$-mixed link diagram, as illustrated in Figure 27. Also, $L$ in the fourth rule stands now for an oriented ${\rm H}$-mixed pseudo link diagram.
@@
Diagrams of the form of Figure 27 have been used in the braid approach to the skein modules of various 3-manifolds. For details cf. [14, 4] and references therein.
@@
Let ${\rm H}\cup K$ be an oriented ${\rm H}$-mixed pseudo link diagram, with ${\rm H}$ being represented in closed braid form. The reduced ${\rm H}$-mixed pseudo bracket polynomial of ${\rm H}\cup K$ is a 5-variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,x^{\pm 1},y\right]$, defined inductively by means of the same rules as for the ${\rm H}$-mixed pseudo bracket polynomial, except for the fifth rule, in which $s_{p,q}$ is replaced by $x^{p}y^{q}$, for $p$ and $q$ pairs of coprime integers with $q$ non negative.
@@
The ${\rm H}$-mixed and the reduced ${\rm H}$-mixed pseudo bracket polynomials are invariant under regular isotopy of oriented ${\rm H}$-mixed pseudo links and are equivalent to the toroidal and the reduced toroidal pseudo bracket polynomials, respectively, for toroidal pseudo links.
@@
We consider the restriction of the theory of planar pseudo links to the subset of ${\rm H}$-mixed pseudo links. Then, by Proposition 1.6, the (reduced) ${\rm H}$-mixed pseudo bracket polynomial is invariant under the classical Reidemeister moves R2 and R3 and under the pseudo Reidemeister moves PR2 and PR3. Regarding now the mixed Reidemeister moves MR2, MR3, MPR3 and mixed R3 (recall Figures 9 and 18), since there is no rule for smoothing mixed crossings, the moves MR2 and the mixed R3 moves are undetectable by the (reduced) ${\rm H}$-mixed pseudo bracket, so it remains invariant. Further, invariance under the moves MR3 and MPR3 follows by the same arguments as in the annular case (recall Figure 14).
@@ 3.4. Jones-type analogues
Let $L$ be an oriented toroidal pseudo link diagram. The normalized toroidal pseudo bracket polynomial is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,s_{p,q}\right]$, defined as:
@@
where $\langle L\rangle_{T^{2}}$ denotes the toroidal pseudo bracket polynomial of $L$ for $H=1-Vd$, and it is an isotopy invariant of toroidal pseudo links and their corresponding pseudo links in the thickened torus.
@@
$P^{r}_{L}(A,V,s_{p,q})\ =\ (-A^{-3})^{w(L)}\,\langle L\rangle^{r}_{T^{2}}$
where $\langle L\rangle^{r}_{T^{2}}$ denotes the reduced toroidal pseudo bracket polynomial of $L$ for $H=1-Vd$, and it is an isotopy invariant of toroidal pseudo links and their corresponding pseudo links in the thickened torus.
Furthermore, for ${\rm H}\cup L$ an oriented ${\rm H}$-mixed pseudo link diagram, the normalized ${\rm H}$-mixed pseudo bracket polynomial is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,s_{p,q}\right]$, defined as:
@@
where $\langle{\rm H}\cup L\rangle$ denotes the ${\rm H}$-mixed pseudo bracket polynomial of ${\rm H}\cup L$ for $H=1-Vd$, and it is an isotopy invariant of ${\rm H}$-mixed pseudo links and their corresponding spatial ${\rm H}$-mixed pseudo links.
@@
$P^{r}_{{\rm H}\cup L}(A,V,s_{p,q})\ =\ (-A^{-3})^{w({\rm H}\cup L)}\,\langle{\rm H}\cup L\rangle^{r},$
where $\langle{\rm H}\cup L\rangle^{r}$ denotes the reduced ${\rm H}$-mixed pseudo bracket polynomial of ${\rm H}\cup L$ for $H=1-Vd$, and it is an isotopy invariant of ${\rm H}$-mixed pseudo links and their corresponding spatial ${\rm H}$-mixed pseudo links.
@@
As in the case of the planar pseudo bracket (Theorem 1.8) and the annular pseudo bracket polynomials (Theorem 2.10), multiplying the toroidal pseudo bracket with the writhe correction factor $\left(-A^{-3}\right)^{w(L)}$ ensures invariance under Reidemeister moves R1, while retaining invariance under regular isotopy, since the writhe remains invariant under regular isotopy. For the invariance under the move PR1, the specialization $H=1-Vd$ must be forced, as argued in [7] and discussed in Subsection 1.4. So, the polynomials $P$ and $P^{r}$ are isotopy invariants of toroidal pseudo links, analogous to the normalized pseudo bracket for planar and annular pseudo links.
@@
Applying the variable substitution $A=t^{-1/4}$ in the polynomials $P_{L}(A,V,s_{p,q})$ and $P^{r}_{L}(A,V,s_{p,q})$ we obtain the equivalent toroidal pseudo Jones polynomials, respectively denoted by $P_{L}(t^{-1/4},V,s_{p,q})$ and $P^{r}_{L}(t^{-1/4},V,s_{p,q})$, which are isotopy invariants of toroidal pseudo links and their corresponding pseudo links in the thickened torus. Similarly, the same variable substitution in the equivalent normalized ${\rm H}$-mixed pseudo bracket polynomials $P_{{\rm H}\cup L}(A,V,s_{p,q})$ and its reduced version $P^{r}_{{\rm H}\cup L}(A,V,s_{p,q})$ gives rise to the ${\rm H}$-mixed pseudo Jones polynomials, denoted $P_{{\rm H}\cup L}(t^{-1/4},V,s_{p,q})$ and $P^{r}_{{\rm H}\cup L}(t^{-1/4},V,s_{p,q})$, isotopy invariants of ${\rm H}$-mixed pseudo links and their spatial analogues.
@@
(a) As in the case of planar and annular pseudo links, the variable $V$ witnesses the presence of pseudo crossings. In the absence of precrossings, the polynomials $P_{L}(A,V,s_{p,q})$, $P^{r}_{L}(A,V,s_{p,q})$, $P_{{\rm H}\cup L}(A,V,s_{p,q})$ and $P^{r}_{{\rm H}\cup L}(A,V,s_{p,q})$ specialize to the analogues of the normalized bracket (or the Jones polynomial for $A=t^{-1/4}$) for links in the thickened torus (for other versions of the bracket and Jones polynomials for classical links in thickened surfaces, we refer to [1, 16]).
(b) The variables $s_{p,q}$ in the toroidal pseudo bracket resp. the ${\rm H}$-mixed pseudo bracket (and variables $x,y$ in the reduced versions) remain unchanged in the normalized polynomials $P_{L}(A,V,s_{p,q})$, $P^{r}_{L}(A,V,s_{p,q})$, $P_{{\rm H}\cup L}(A,V,s_{p,q})$ and $P^{r}_{{\rm H}\cup L}(A,V,s_{p,q})$, and they distinguish essential toroidal pseudo links from non-essential ones (i.e. ones that can be isotoped in an inclusion disc). Clearly, for non-essential toroidal pseudo links, the above normalized polynomials specialize to the corresponding planar one, $P_{L}(A,V)$.
(c) The variables $s_{p,q}$ in the normalized polynomials $P_{L}(A,V,s_{p,q})$ and $P_{{\rm H}\cup L}(A,V,s_{p,q})$ (and variables $x,y$ in the reduced versions $P^{r}_{L}(A,V,s_{p,q})$ and $P^{r}_{{\rm H}\cup L}(A,V,s_{p,q})$) can also distinguish essential toroidal pseudo links from annular ones.

### v2

@@ 3.1. Preliminaries
Figure 17. The lift of a toroidal pseudo link to a pseudo link in $T^{2}\times I$.
@@
Theorem 3.2 ([7]).
Two oriented $\rm H$}-mixed pseudo links in $S^{3}$ are isotopic if and only if any two oriented $\rm H$}-mixed pseudo link diagrams of theirs differ by planar isotopies, a finite sequence of the classical and the pseudo Reidemeister moves for the moving parts of the mixed pseudo links (recall Figure 2), and moves that involve the fixed and the moving parts, called mixed Reidemeister moves, comprising the moves MR2, MR3, and MPR3, as exemplified in Figure 9 and the mixed R3 moves, as exemplified in Figure 19.
@@ 3.2. The toroidal pseudo bracket polynomial
A $(p,q)$-torus knot, for $p$ and $q$ relatively prime integers, is an essential simple closed curve in the torus $T^{2}$ winding $p$ times around the longitude and $q$ times along the meridian of the torus. Similarly, a $(p^{\prime},q^{\prime})$-torus link or a $(p,q)$-torus link on $l$ components is a set of $l$ parallel copies of a $(p,q)$-torus knot, where $p^{\prime}=l\cdot p$ and $q^{\prime}=l\cdot q$. In our definition of torus knots and links, we include the pairs $(p,0)$ and $(0,q)$.
@@
A $(p,q)$-torus knot, considered as a classical knot in $S^{3}$, is equivalent a $(q,p)$-torus knot. However, as essential curves embedded in the torus, these two knots are distinct in $T^{2}$. Yet, note that in $T^{2}$ a $(p,q)$-torus knot is the same as a (p,q)(-p,-q)-torus knot, and similarly, a (p,q)(p,-q)-torus knot is the same as a (p,q)(-p,q)-torus knot.
All torus knots and links along with the standard null-homotopic unknot constitute initial conditions for the pseudo bracket polynomial of toroidal pseudo link diagrams. So, we assign a specific variable $s_{p,q}$ to each $(p,q)$-torus knot and we define the following:
@@
Let $K$ be an oriented toroidal pseudo link diagram. The toroidal pseudo bracket polynomial of $K$, denoted $\langle K\rangle_{T^{2}}$ or simply $\langle K\rangle$, is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,s_{p,q}\right]$, for $p$ and $q$ pairs of coprime integers with $q$ non negative, defined inductively by means of the following rules, or $(p,q)\in\{(1,0),(0,1)\}$, where a punctured plane in the 5th rule represents the torus:
@@
Let $K$ be an oriented toroidal pseudo link. The reduced toroidal pseudo bracket polynomial of $K$, denoted $\langle K\rangle^{r}_{T^{2}}$ or simply $\langle K\rangle^{r}$, is a 5-variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,x^{\pm 1},y\right]$, defined inductively by means of the same rules as for the polynomial $\langle K\rangle_{T^{2}}(A,V,H,s_{p,q})$, except for the fifth rule, in which $s_{p,q}$ is replaced by $x^{p}y^{q}$, for $p$ and $q$ pairs of coprime integers with $q$ non negative, or $(p,q)\in\{(1,0),(0,1)\}$. We further define the universal toroidal pseudo bracket polynomial by substituting in the fifth rule $s_{p,q}^{k}$ by $s_{p,q,k}$.
Remark 3.6.
Since the precrossings are smoothened, the universal toroidal pseudo bracket polynomial recovers the Kauffman bracket skein module of toroidal pseudo links. Cf. [1] and references therein.
Note that the last rule can only be applied to a toroidal state diagram that contains only a standard null-homotopic unknot. We also point out that two torus knot components of a toroidal pseudo link diagram, which have different slopes, will necessarily form crossings, which will be smoothened by the skein relations of Definition 3.4 (and Definition 3.5 for the reduced case). So, a state diagram of the toroidal pseudo bracket polynomial can only contain a single torus knot or link. Therefore, together with the comment above Definition 3.4 we have that the toroidal pseudo bracket polynomial and the reduced one are well defined.
@@
The (universal) toroidal pseudo bracket polynomial and the reduced toroidal pseudo bracket polynomial are invariant under regular isotopy of oriented toroidal pseudo links.
@@
In [7] we discuss inclusion relations of the pseudo knot theories of the various supporting manifolds involved. Based on these inclusion relations we now discuss relations of the corresponding pseudo bracket polynomials.
@@
Figure 20. Inclusion relations: a disc in the torus and an annulus in the torus.
@@
Figure 21. Inclusion relation of the thickened torus in a 3-ball.
@@
Figure 22. Inclusion relation of the thickened torus in the solid torus.
@@
Remark 3.9.
Consider a torus link consisting of two different torus knot components, say a $(1,0)$-torus knot and a $(0,1)$-torus knot which intersect in a precrossing. Assigning orientations to the two components we observe that reversing the orientation of one component will result in different bracket polynomials, due to the fact that the roles of $V$ and $H$ in the third rule of the toroidal pseudo bracket are not interchangeable. The situation is illustrated in Figure 24. The left-hand side toroidal pseudo link, $L_{1}$, is assigned the polynomial:
@@
The pseudo trefoil on the top right side of the figure, denoted $K_{r}$, is not of annular type, so when projecting on the annulus, a virtual crossing will appear, as illustrated in Figure 25 bottom right, which is neither a real crossing nor a precrossing (cf. [16]). The use of virtual crossings in this context comes in handy (only) for illustration purposes, as discussed in [7], in order to represent toroidal diagrams in the annulus or in the once punctured disc. We stress that the isotopy moves of virtual knot theory do not apply here. The skein tree analysis of $K_{r}$ is depicted in Figure 26.
@@ 3.3. The $\rm H$}-mixed pseudo bracket polynomial
We now consider the torus links of type $(kp,kq)$, which form the states for the toroidal pseudo bracket polynomial of Definition 3.4, and which we want to correspond to $\rm H$}-mixed links. Figure 27 illustrates a $(3,2)$-torus knot represented in two ways as an $\rm H$}-mixed link. The left-hand illustration of the figure uses the common representation of $\rm H$}, while in the right-hand illustration $\rm H$} is represented in closed braid form. As we observe, in the left-hand illustration appear three moving crossings, which should be smoothened according to the bracket rules, a situation we can avoid using the representation of the $(3,2)$-torus knot in the right-hand illustration.
@@
We are now ready to translate the toroidal pseudo bracket polynomial in the $\rm H$}-mixed setting:
@@
Let $K$ be a toroidal oriented pseudo link diagram and let ${\rm H}\cup K$ the corresponding $\rm H$}-mixed pseudo link diagram, with $\rm H$} being represented in closed braid form. The $\rm H$}-mixed pseudo bracket polynomial of ${\rm H}\cup K$ is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,s_{p,q}\right]$, defined by means of the same inductive rules as the ones for the toroidal mixed pseudo bracket polynomial, except for the $(kp,kq)$-torus link diagram in the fifth rule of Definition 3.4, which is substituted by the corresponding $\rm H$}-mixed link diagram, as illustrated in Figure 28. Also, $L$ in the fourth rule stands now for an oriented $\rm H$}-mixed pseudo link diagram. We further define the universal $\rm H$}-mixed pseudo bracket polynomial by substituting in the fifth rule $s_{p,q}^{k}$ by $s_{p,q,k}$.
@@
Diagrams of the form of Figure 28 have been used in the braid approach to the skein modules of various 3-manifolds. Note that these diagrams comprise only a subset of the complete theory of $\rm H$}-mixed links, where different torus links may be present in a state diagram. For details cf. [3] and references therein.
@@
Let ${\rm H}\cup K$ be an oriented $\rm H$}-mixed pseudo link diagram, with $\rm H$} being represented in closed braid form. The reduced $\rm H$}-mixed pseudo bracket polynomial of ${\rm H}\cup K$ is a 5-variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,H,x^{\pm 1},y\right]$, defined inductively by means of the same rules as for the $\rm H$}-mixed pseudo bracket polynomial, except for the fifth rule, in which $s_{p,q}$ is replaced by $x^{p}y^{q}$, for $p$ and $q$ pairs of coprime integers with $q$ non negative, or $(p,q)\in\{(1,0),(0,1)\}$.
@@
The (universal) $\rm H$}-mixed and the reduced $\rm H$}-mixed pseudo bracket polynomials are invariant under regular isotopy of oriented $\rm H$}-mixed pseudo links and are equivalent to the (universal) toroidal and the reduced toroidal pseudo bracket polynomials, respectively, for toroidal pseudo links.
@@
We consider the restriction of the theory of planar pseudo links to the subset of $\rm H$}-mixed pseudo links. Then, by Proposition 1.6, the (reduced) $\rm H$}-mixed pseudo bracket polynomial is invariant under the classical Reidemeister moves R2 and R3 and under the pseudo Reidemeister moves PR2 and PR3. Regarding now the mixed Reidemeister moves MR2, MR3, MPR3 and mixed R3 (recall Figures 9 and 19), since there is no rule for smoothing mixed crossings, the moves MR2 and the mixed R3 moves are undetectable by the (reduced) $\rm H$}-mixed pseudo bracket, so it remains invariant. Further, invariance under the moves MR3 and MPR3 follows by the same arguments as in the annular case (recall Figure 15).
@@ 3.4. Jones-type analogues
Let $L$ be an oriented toroidal pseudo link diagram. The (universal) normalized toroidal pseudo bracket polynomial is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,s_{p,q}\right]$ $\left(\text{resp.}\,\mathbb{Z}\left[A^{\pm 1},V,s_{p,q,k}\right]\right)$, defined as:
@@
$\left(\text{resp.}\,P_{L}(A,V,s_{p,q,k})\right)$ where $\langle L\rangle_{T^{2}}$ denotes the (universal) toroidal pseudo bracket polynomial of $L$ for H=1VdH=1-Vd, and it is an isotopy invariant of toroidal pseudo links and their corresponding pseudo links in the thickened torus.
@@
$P^{r}_{L}(A,V,x,y)\ =\ (-A^{-3})^{w(L)}\,\langle L\rangle^{r}_{T^{2}}$
where $\langle L\rangle^{r}_{T^{2}}$ denotes the reduced toroidal pseudo bracket polynomial of $L$ for H=1VdH=1-Vd, and it is an isotopy invariant of toroidal pseudo links and their corresponding pseudo links in the thickened torus.
Furthermore, for ${\rm H}\cup L$ an oriented $\rm H$}-mixed pseudo link diagram corresponding to the toroidal pseudo link diagram $L$, the (universal) normalized $\rm H$}-mixed pseudo bracket polynomial is an infinite variable Laurent polynomial in the ring $\mathbb{Z}\left[A^{\pm 1},V,s_{p,q}\right]$, defined as:
@@
$\left(\text{resp. }P_{{\rm H}\cup L}(A,V,s_{p,q,k})\text{ in the ring }\mathbb{Z}\left[A^{\pm 1},V,s_{p,q,k}\right]\right)$ where $\langle{\rm H}\cup L\rangle$ denotes the (universal) $\rm H$}-mixed pseudo bracket polynomial of ${\rm H}\cup L$ for H=1VdH=1-Vd, and it is an isotopy invariant of $\rm H$}-mixed pseudo links and their corresponding spatial $\rm H$}-mixed pseudo links.
@@
$P^{r}_{{\rm H}\cup L}(A,V,x,y)\ =\ (-A^{-3})^{w({\rm H}\cup L)}\,\langle{\rm H}\cup L\rangle^{r},$
where $\langle{\rm H}\cup L\rangle^{r}$ denotes the reduced $\rm H$}-mixed pseudo bracket polynomial of ${\rm H}\cup L$ for H=1VdH=1-Vd, and it is an isotopy invariant of $\rm H$}-mixed pseudo links and their corresponding spatial $\rm H$}-mixed pseudo links.
@@
As in the case of the planar pseudo bracket (Theorem 1.8) and the annular pseudo bracket polynomials (Theorem 2.10), multiplying the (universal) toroidal pseudo bracket with the writhe correction factor $\left(-A^{-3}\right)^{w(L)}$ ensures invariance under Reidemeister moves R1, while retaining invariance under regular isotopy, since the writhe remains invariant under regular isotopy. For the invariance under the move PR1, the specialization H=1VdH=1-Vd must be forced, as argued in [9] and discussed in Subsection 1.4. So, the polynomials $P$ and $P^{r}$ are isotopy invariants of toroidal pseudo links, analogous to the normalized pseudo bracket for planar and annular pseudo links.
@@
Applying the variable substitution $A=t^{-1/4}$ in the polynomials $P_{L}(A,V,s_{p,q,k})$, $P_{L}(A,V,s_{p,q})$ and $P^{r}_{L}(A,V,x,y)$ we obtain the equivalent (universal) toroidal pseudo Jones polynomials, respectively denoted by $P_{L}(t^{-1/4},V,s_{p,q,k})$, $P_{L}(t^{-1/4},V,s_{p,q})$ and $P^{r}_{L}(t^{-1/4},V,x,y)$, which are isotopy invariants of toroidal pseudo links and their corresponding pseudo links in the thickened torus. Similarly, the same variable substitution in the equivalent normalized $\rm H$}-mixed pseudo bracket polynomials $P_{{\rm H}\cup L}(A,V,s_{p,q,k})$, $P_{{\rm H}\cup L}(A,V,s_{p,q})$ and the reduced version $P^{r}_{{\rm H}\cup L}(A,V,x,y)$ gives rise to the $\rm H$}-mixed pseudo Jones polynomials, denoted $P_{{\rm H}\cup L}(t^{-1/4},V,s_{p,q,k})$, $P_{{\rm H}\cup L}(t^{-1/4},V,s_{p,q})$ and $P^{r}_{{\rm H}\cup L}(t^{-1/4},V,x,y)$, isotopy invariants of $\rm H$}-mixed pseudo links and their spatial analogues.
@@
(a) As in the case of planar and annular pseudo links, the variable $V$ witnesses the presence of pseudo crossings. In the absence of precrossings, the polynomials $P_{L}(A,V,s_{p,q,k})$, $P_{L}(A,V,s_{p,q})$, $P^{r}_{L}(A,V,x,y)$, $P_{{\rm H}\cup L}(A,V,s_{p,q,k})$, $P_{{\rm H}\cup L}(A,V,s_{p,q})$ and $P^{r}_{{\rm H}\cup L}(A,V,x,y)$ specialize to the analogues of the normalized bracket (or the Jones polynomial for $A=t^{-1/4}$) for links in the thickened torus (for other versions of the bracket and Jones polynomials for classical links in thickened surfaces, we refer to [2, 21]).
(b) The variables $s_{p,q}$ (resp. $s_{p,q,k}$) in the (universal) toroidal pseudo bracket resp. the (universal) $\rm H$}-mixed pseudo bracket (and variables $x,y$ in the reduced versions) remain unchanged in the normalized polynomials $P_{L}(A,V,s_{p,q,k})$, $P_{L}(A,V,s_{p,q})$, $P^{r}_{L}(A,V,x,y)$, $P_{{\rm H}\cup L}(A,V,s_{p,q,k})$, $P_{{\rm H}\cup L}(A,V,s_{p,q})$ and $P^{r}_{{\rm H}\cup L}(A,V,x,y)$, and they distinguish essential toroidal pseudo links from non-essential ones (i.e. ones that can be isotoped in an inclusion disc). Clearly, for non-essential toroidal pseudo links, the above normalized polynomials specialize to the corresponding planar one, $P_{L}(A,V)$.
(c) The variables $s_{p,q}$ (resp. $s_{p,q,k}$) in the normalized (universal) polynomials $P_{L}(A,V,s_{p,q,k})$, $P_{L}(A,V,s_{p,q})$ and $P_{{\rm H}\cup L}(A,V,s_{p,q,k})$, $P_{{\rm H}\cup L}(A,V,s_{p,q})$ (and variables $x,y$ in the reduced versions $P^{r}_{L}(A,V,x,y)$ and $P^{r}_{{\rm H}\cup L}(A,V,x,y)$) can also distinguish essential toroidal pseudo links from annular ones.